Is this the right way do do the problem: 16-3(8-3)^2/5
13(5)2power/5
13(25)/5
325/5
65

Answer by praseenakos@yahoo.com(507)


16-3(8-3)^2/5
16-3(5)^2/5
16-3.5= 16-15=1 is the correct way of doing the sum